# Data for Mathematical Copilots: Better Ways of Presenting Proofs for Machine Learning

### v1

@@ 4.1 Mapping Mathematical Workflows to Data
  • To define a Floer-theoretic invariant, we construct a chain complex whose boundary map counts some pseudo-holomorphic curves with certain Lagrangian boundary conditions in some symplectic manifold. To show these counts are finite, one has to prove transversality results for the moduli spaces using difficult methods from PDEs, then obtain a formula for the dimension of the moduli spaces and compactify the moduli spaces. The next step is showing the boundary map squares to zero, in which case we have a chan complex, or sometimes we end up with a more complicated algebraic structure, such as an $A^{\infty}$-module or a differential graded algebra. One often has to use special coefficient systems, such as a Novikov ring, and deal with bubbling phenomena. There are many choices that go into the construction of our chain complexes, so one has to prove independence of these choices up to chain homotopy equivalence, which often relies on continuation maps and pseudo-holomorphic polygon counts. One then extracts more tractable invariants from the chain homotopy type using algebra, such as taking homology. Different invariants are often related by spectral sequences, defined using filtrations of the chain complexes. A similar workflow exists for gauge-theoretic invariants. More sophisticated invariants can be obtained by exploiting certain symmetries, such as $\mathbb{Z}_{2}$ or $\text{Pin}(2)$. In another popular direction, one can sometimes construct a homotopy type from chain complexes; see (Lipshitz and Sarkar, 2018). Certain constructions from algebraic topology can then be invoked to provide additional algebraic structure, such as Steenrod operations. We can hence represent the high-level workflow for defining a Floer-theoretic topological invariant as follows:
@@
    Certain knot invariants are easy to define, but no algorithm is known to compute them, such as the unknotting number or the 4-ball genus, and topologists use computable invariants, many of them arising from knot Floer homology, to give lower bounds on these. The paper (Davies et al., 2021) describes a workflow for using supervised learning to find correlations between various invariants, which has led to an inequality relating the knot signature and hyperbolic invariants. Upper bounds can be obtained by performing certain moves on knot diagrams, possibly with the help of reinforcement learning or Bayesian optimization; see (Gukov et al., 2023) and (Applebaum et al., 2024). For example, one could take (4-ball genus, lower bound from Rasmussen $s$-invariant, torus knots), which leads to a solution of Milnor’s conjecture on the 4-ball genus of torus knots (Rasmussen, 2010).
@@
  • The classification of smooth 4-manifolds is still wide open. There exist 4-manifolds that are homeomorphic but not diffeomorphic, which are called exotic pairs. To construct these, experts use methods from algebraic and symplectic geometry, such as blow-ups, fiber sums when given Lefschetz fibrations, and other gluings, typically along 3-tori, such as knot surgery. To show two 4-manifolds are homeomorphic, one shows they have isomorphic intersection forms and invokes Freedman’s theorem (Freedman, 1982), for which the fundamental group has to satisfy certain restrictions (e.g., being trivial). The intersection form and a presentation of the fundamental group can be read off a Kirby diagram. This fundamental group computation is often the most difficult step, and note that any finitely presented group can arise as the fundamental group of a closed 4-manifold. Furthermore, there is no algorithm to decide whether a finitely presented group is trivial. So, some questions in low-dimensional topology are beyond the reach of computers. As mentioned above, one can use Kirby calculus to show if two 4-manifolds are diffeomorphic. Swenton’s Knot-Like Objects software [^13] https://community.middlebury.edu/~mathanimations/klo/ is capable of performing Kirby moves. To show they are not diffeomorphic, one usually distinguishes them using the gauge-theoretic Seiberg–Witten invariants. No algorithm is known to compute these, but they can be computed in some instances using Taubes’ non-vanishing result for symplectic 4-manifolds (Taubes, 1994), together with various glueing results, such as Fintushel and Stern’s knot surgery formula (Fintushel and Stern, 1998). A schematic of a typical workflow in smooth 4-manifold topology is as follows:
@@
    choose method of construction $\rightarrow$ construct pair of smooth 4-manifolds $\rightarrow$ find their Kirby diagrams $\rightarrow$ show they are simply-connected $\rightarrow$ show they have isomorphic intersection forms (hence homeomorphic by Freedman) $\rightarrow$ try to show they are diffeomorphic using Kirby calculus; if this fails $\rightarrow$ compute Seiberg–Witten invariants to show not diffeomorphic.
@@
3-manifold topology has a completely different flavour. Here, methods from geometric group theory, hyperbolic geometry, and combinatorial topology dominate. These combinatorial methods include triangulations and normal surface theory. Geometric group theory studies groups using their actions on metric spaces (e.g., on the Cayley graph). Also note Perelman’s proof of the 3-dimensional Poincaré conjecture using the Ricci flow, rooted purely in geometric analysis; see (Morgan and Tian, 2007). Well-developed computer packages exist to aid 3-manifold topologists, including the already mentioned SnapPy for hyperbolic 3-manifolds and Regina (Burton, 2004) for triangulations. We do not provide specific workflows in 3-manifold topology due to the diversity of techniques used.
@@
The above-mentioned question then becomes to find a sequence of deformations of the embedded circle such that its projection has no crossings. It is known (Kauffman and Lambropoulou, 2012) that for some data representations of the projected knot (such as Dowker–Thistlethwaite codes (Dowker and Thistlethwaite, 1983) with Reidemeister moves as allowed deformations (Reidemeister, 1927)), there exist examples where the number of crossings needs to be increased before all crossings can be removed. In another representation (grid representation with Dynnikov moves as allowed deformations), the number of crossings is monotonically decreasing (Dynnikov, 2006), but both the representation of the knot as a grid diagram and the carrying out of the simplification steps are more complicated. These distinct representations of the same mathematical object lend themselves to different foundational models (Gukov et al., 2021; Kauffman et al., 2022); for example, braid word representations of knot projections are closer to natural language, whereas other representations are closer to vision tasks or graphs. The different representations also inform the type of algorithm that needs to be used. If the simplification is non-monotonic, a local search can get stuck in local minima.
@@
One common activity in mathematics, when confronted with a new statement of unknown truth, a conjecture, is to engage in the search for proof – or a counterexample. This comes with its own workflows. The search for counterexamples to conjectures is particularly useful for hard problems with a truth certificate that can be verified in polynomial time, such as NP-hard or NP-complete problems. The idea is to cast the search problem into a Markov Decision Problem (MDP) whose terminal states are counterexamples, and attempt to solve the MDP using data science techniques. In the past, deep reinforcement learning has proven to be a powerful tool. If the RL agent finds a solution to the MDP, its episodic rollouts provide truth certificates for the counterexample, thus establishing a verifiably correct proof by counter-example, see e.g. (Gukov et al., 2024) for a recent summary of these ideas and (Gukov et al., 2021; Wagner, 2021; Gukov et al., 2023; Charton et al., 2024) for some recent concrete applications.
@@ 4.2 Data Collection in Real Environments
To gather representative data on mathematical workflow steps, as well as other metamathematical items, such as proof heuristics, limitations of certain proof techniques, etc., it may be desirable to observe and absorb the full process of producing mathematics, including all data that is not. For an example of real-time narration of thought processes that arise when doing competitive problem-solving, see co-author Tim Gowers’ series of YouTube videos. [^14] https://www.youtube.com/@TimothyGowers0/videos
Unfortunately, many of these intermediate process steps become evident only implicitly from data sources. For example, at the elementary level, various workflows to solve various limits,
@@
which require repeated applications of known theorems, such as de l’Hôpital’s theorem, or related ones to resolve indeterminate cases, are distilled only by solving a large number of exercises; no explicit annotations for these workflows exist. On a more advanced level, these are often conveyed in blogs, talks, and oral discussions between mathematicians at conferences – and not in the typical data sources used to train LLMs.
@@
In principle, one could implement longitudinal studies that track research projects from inception to completion. However, such data collection would require classifying, structuring, and recording different tasks throughout the mathematician’s daily work. The apparent challenge is to scale such approaches while keeping the overhead as small as possible for the researcher whose work is tracked.
@@
While such approaches are welcome to be tested, as a first alternative step, we recommend transcribing from online sources such as lectures, panel discussions, seminars, vlogs, etc. While such data only requires minimal need for postprocessing and is often already collected in the form of videos, it typically covers broader mathematical aspects than textbooks. However, while humans can frequently generalize from a single instance of a certain workflow, LLMs often require a dataset that contains sufficiently many examples. While data collection effort will elicit higher-quality mathematical data, the time lag between the time when new research discoveries are made and when these are represented in lectures can be significant. On the other hand, at venues such as conferences, a lot of information about mathematics is exchanged, but not recorded. While it would technically be possible to record a portion of the conversations that happen at a conference in a privacy-preserving manner, and this would be a great source of data, as it one would consistently have data that is at the forefront of research (unlike the mentioned panel discussions, where there is often a time lag), it would require a paradigm shift from conference participants to accept such data collecting measures. We urge further conversations about privacy-preserving ways of curating richer datasets on mathematical workflows, that minimally impinge on mathematicians’ naturalistic practices.

### v2

@@ 4.1 Mapping Mathematical Workflows to Data
  • To define a Floer-theoretic invariant, we construct a chain complex whose boundary map counts some pseudo-holomorphic curves with certain Lagrangian boundary conditions in some symplectic manifold. To show these counts are finite, one has to prove transversality results for the moduli spaces using difficult methods from PDEs, then obtain a formula for the dimension of the moduli spaces and compactify the moduli spaces. The next step is showing the boundary map squares to zero, in which case we have a chain complex, or sometimes we end up with a more complicated algebraic structure, such as an $A^{\infty}$-module or a differential graded algebra. One often has to use special coefficient systems, such as a Novikov ring, and deal with bubbling phenomena. There are many choices that go into the construction of our chain complexes, so one has to prove independence of these choices up to chain homotopy equivalence, which often relies on continuation maps and pseudo-holomorphic polygon counts. One then extracts more tractable invariants from the chain homotopy type using algebra, such as taking homology. Different invariants are often related by spectral sequences, defined using filtrations of the chain complexes. A similar workflow exists for gauge-theoretic invariants. More sophisticated invariants can be obtained by exploiting certain symmetries, such as $\mathbb{Z}_{2}$ or $\text{Pin}(2)$. In another popular direction, one can sometimes construct a homotopy type from chain complexes; see (LipshitzSarkar:2018). Certain constructions from algebraic topology can then be invoked to provide additional algebraic structure, such as Steenrod operations. We can hence represent the high-level workflow for defining a Floer-theoretic topological invariant as follows:
@@
    Certain knot invariants are easy to define, but no algorithm is known to compute them, such as the unknotting number or the 4-ball genus, and topologists use computable invariants, many of them arising from knot Floer homology, to give lower bounds on these. The paper (davies2021advancing) describes a workflow for using supervised learning to find correlations between various invariants, which has led to an inequality relating the knot signature and hyperbolic invariants. Upper bounds can be obtained by performing certain moves on knot diagrams, possibly with the help of reinforcement learning or Bayesian optimization; see (Gukov:2023SearchingRibbons) and (unknotting2024). For example, one could take (4-ball genus, lower bound from Rasmussen $s$-invariant, torus knots), which leads to a solution of Milnor’s conjecture on the 4-ball genus of torus knots (Rasmussen:2010slice).
@@
  • The classification of smooth 4-manifolds is still wide open. There exist 4-manifolds that are homeomorphic but not diffeomorphic, which are called exotic pairs. To construct these, experts use methods from algebraic and symplectic geometry, such as blow-ups, fiber sums when given Lefschetz fibrations, and other gluings, typically along 3-tori, such as knot surgery. To show two 4-manifolds are homeomorphic, one shows they have isomorphic intersection forms and invokes Freedman’s theorem (Freedman:1982four-manifolds), for which the fundamental group has to satisfy certain restrictions (e.g., being trivial). The intersection form and a presentation of the fundamental group can be read off a Kirby diagram. This fundamental group computation is often the most difficult step, and note that any finitely presented group can arise as the fundamental group of a closed 4-manifold. Furthermore, there is no algorithm to decide whether a finitely presented group is trivial. So, some questions in low-dimensional topology are beyond the reach of computers. As mentioned above, one can use Kirby calculus to show if two 4-manifolds are diffeomorphic. Swenton’s Knot-Like Objects software [^14] https://community.middlebury.edu/~mathanimations/klo/ is capable of performing Kirby moves. To show they are not diffeomorphic, one usually distinguishes them using the gauge-theoretic Seiberg–Witten invariants. No algorithm is known to compute these, but they can be computed in some instances using Taubes’ non-vanishing result for symplectic 4-manifolds (Taubes:1994), together with various glueing results, such as Fintushel and Stern’s knot surgery formula (FintushelStern:1998). A schematic of a typical such workflow in smooth 4-manifold topology is as follows:
@@
    pair of smooth 4-manifolds $\rightarrow$ find their Kirby diagrams $\rightarrow$ show they are simply-connected $\rightarrow$ show they have isomorphic intersection forms (hence homeomorphic by Freedman) $\rightarrow$ try to show they are diffeomorphic using Kirby calculus; if this fails $\rightarrow$ compute Seiberg–Witten invariants to show not diffeomorphic.
@@
3-manifold topology has a completely different flavour. Here, methods from geometric group theory, hyperbolic geometry, and combinatorial topology dominate. These combinatorial methods include triangulations and normal surface theory. Geometric group theory studies groups using their actions on metric spaces (e.g., on the Cayley graph). Also note Perelman’s proof of the 3-dimensional Poincaré conjecture using the Ricci flow, rooted purely in geometric analysis; see (MorganTian:2007). Well-developed computer packages exist to aid 3-manifold topologists, including the already mentioned SnapPy for hyperbolic 3-manifolds and Regina (Burton:2004Regina) for triangulations. We do not provide specific workflows in 3-manifold topology due to the diversity of techniques used.
Having discussed differential equations and low dimensional topology, we now turn to the Langlands philosophy, a far-reaching series of conjectures relating the continuous to the discrete. The story starts with Gauss’ law of quadratic reciprocity, a striking result explaining a rather counterintuitive relationship between two apparently unrelated questions; given two odd primes $p$ and $q$, there is an unexpected symmetry relating the questions of whether $p$ is a square modulo $q$, and whether $q$ is a square modulo $p$. Gauss’ theorem can be thought of as one of the first nontrivial results in arithmetic. In the 1800s Eisenstein extended the result to higher powers, and ultimately the results were recast as class field theory, one of the highlights of 19th century mathematics. The advent of cohomology (a tool initially introduced to study questions in topology) after the war clarified things further, and by the 1950s the main proofs (now in the form of Artin reciprocity) were well-understood, although they even now remain frustratingly long. For example the 1968 book by Cassels and Froehlich cassels1976algebraic which contains the proofs of the main theorems is over 300 pages long, and the proofs have not become substantially shorter since.
The bombshell dropped by Langlands in the 1960s was to observe that class field theory itself was merely the 1-dimensional case of a far more profound nonabelian reciprocity law relating objects from harmonic analysis (the continuous) and objects from arithmetic (the discrete). In short, there seemed to be a collection of analytic objects (“automorphic representations”) each of which was related to a corresponding algebraic object (a “Galois representation”). Rest assured that the details of the (highly technical) definitions are irrelevant to the story being told here; the key point is that automorphic representations are objects coming from harmonic analysis, and Galois representations are objects coming from arithmetic, and these are (on the face of it) two very different areas of mathematics.
Whilst the initial idea was compelling, it took some time before the depths of Langlands’ insights could be turned into rigorous mathematical statements, with early attempts involving data whose existence was only conjectural (for example the global Langlands group of a number field, an object whose definition is still unclear). On the other hand, a wealth of evidence accrued very quickly, making it absolutely clear that beneath the “philosophy” there were some beautiful mathematical statements waiting to be proved. In the 60 years since Langlands’ ideas were first enunciated, there have been generalizations to other domains of mathematics; there are local and global Langlands conjectures, a $p$-adic Langlands philosophy, a mod $p$ philosophy (both of which are still wide-open and whose statements are still not completely understood), and geometric Langlands conjectures, where our understanding is much better and there is even a recent announcement of a proof in many cases, which was picked up by the media, although the papers have not yet been published.
Let us stick to the classical local and global conjectures, which already raise plenty of questions. What do the workflows look like in this area?
  • There is the fight to understand the question. Part of the subtlety of this part of mathematics is that whilst there is a general idea that certain automorphic representations (objects from analysis) should be in correspondence with certain Galois representations (objects from arithmetic), it is still not clear how to uniquely pin down such a correspondence (for general reductive groups over number fields, say). One can write down a list of criteria which the correspondence should satisfy, but what mathematicians really seek is some kind of constructive or “canonical” construction of an object on one side of the conjectured correspondence, given an object on the other side. Without this definition, it is impossible to formally state the theorems we would like to prove. In particular, the main goal here is not to compute a value or to prove a theorem, it is to construct a function.
  • However there are plenty of instances of rigorous mathematical conjectures which we can write down and for which we have ample evidence in the form of both numerical examples and partial results. Here is a conjecture which can be made completely precise. Let $n$ be a positive integer. Given an algebraic automorphic representation for the group $GL(n)$ over a number field, one expects to be able to associate to it a compatible family of $n$-dimensional $p$-adic Galois representations of this number field, and conversely given the family of Galois representations one expects the existence of an automorphic representation. This statement is fully proved in the case $n=1$ by class field theory. What little we do know for $n>1$ is for the most part restricted to totally real and CM number fields (which together form 0 percent of all number fields, one fact indicating the depths of our ignorance) and what litle we do know has been very hard-won. The key difficulty is to find some kind of bridge taking us between the discrete and the continuous, and there are limited tools available. Certain automorphic representations can be studied using differential geometry. Serre’s work linked differential geometry and algebraic geometry in the 1950s; Grothendieck’s revolution of algebraic geometry in the 1960s gave links to arithmetic, and thus provided the first viable bridge in the case of $n=2$. Algebraic geometry was the source of the first general constructions of elliptic curves from modular forms, provided by Eichler and Shimura at the end of that decade. This work was simultaneously a milestone achievement and a miniscule special case of the Langlands program for $GL(2)$ over the rationals. We had to wait for 25 years before Wiles and Taylor managed to go from an elliptic curve back to a modular form; Fermat’s Last Theorem was a consequence. In particular a major problem in number theory was resolved by proving an extremely small part of Langlands’ conjectures. One of the major difficulties here is that the arguments need nontrivial inputs from arithmetic, algebra, geometry and analysis, and would take thousands of pages to explain if written out in full. These thousands of pages establish just a very small special case of the Langlands program for $GL(2)$ over the rationals. In some sense it does not bode well that the $GL(1)$ proofs were hundreds of pages long and even a small fragment of the $GL(2)$ proofs are thousands of pages. Within these pages there are of course several profound ideas; many more seem to be needed. There are few living humans who have read and completely understood the full proofs of the modularity theorem linking elliptic curves and modular forms; on the other hand it is certainly possible to understand a high-level overview, and most people take on parts of the argument as “black boxes” where the statements of the results are comprehensible but the proofs might involve techniques which are not part of their expertise. In particular the workflow here does not involve understanding all the details of everything; communication between mathematicians with different expertise is key here, as well as taking results from other areas of mathematics on trust.
  • The success of Wiles’ program showed in particular that algebraic geometry gave a viable route to cross the chasm between the discrete and the continuous and link objects on both sides. However the techniques will only apply when certain quite restrictive conditions hold. Another tool in the area is “Langlands functoriality”, a tool for producing new analytic objects from old ones based on predictions on the arithmetic side of the picture. If the Langlands conjectures are true then basic constructions in arithmetic should have analogues on the analytic side, and one can independently try and make those constructions. Another motivation for making such constructions is that one might start with an analytic object for which the algebro-geometric tools do not apply, and then attempt to massage it until they do apply. Many of the results known for $GL(n)$ for $n>2$ are proved in this way; the group $GL(n)$ does “not admit Shimura varieties” for $n>2$, meaning that algebraic geometry has limited applicability; however if one can use analytic means to move from an automorphic representation on $GL(n)$ to an automorphic representation on a unitary group (where the methods do work) then this enables more analytic objects to “cross the bridge” to the algebraic side. Results in Langlands functoriality thus have the dual role of adding weight to the conjectures (by proving consequences of them) and also enabling us to prove more cases of the conjectures. Here the work is Fourier-analytic and group-theoretic, with less reliance on algebraic and arithmetic geometry.
  • As is often the case in mathematics, people attempt to understand the questions in this area by generalizing them. The Galois representations which the Langlands philosophy are $p$-adic in nature, and can move freely in $p$-adic families. This raises the question of whether one can $p$-adically deform automorphic representations; initially this sounds ridiculous because these objects are inherently complex-analytic. However $p$-adic analytic variants of the theory have now been established, and this gives us new freedom, unlocking new tools to establish more cases of the classical program; the key observation here is that $p$-adic analysis is far more closely related to arithmetic than classical analysis. What is happening in this area is a mixture of insights gleaned from analogy (“what if there is a more arithmetic version of analysis?”) turn into an actual theory (the existence of $p$-adic automorphic representations) but then must be backed up by very technical work showing that the theory can actually be used to solve new cases of the problem at hand. Work in this area thus needs to blend heuristic ideas and technical calculations.
@@
The above-mentioned question then becomes to find a sequence of deformations of the embedded circle such that its projection has no crossings. It is known (Kauffman:2012Hard) that for some data representations of the projected knot (such as Dowker–Thistlethwaite codes (Dowker:1982Classification) with Reidemeister moves as allowed deformations (Reidemeister:1927Abhandlungen)), there exist examples where the number of crossings needs to be increased before all crossings can be removed. In another representation (grid representation with Dynnikov moves as allowed deformations), the number of crossings is monotonically decreasing (Dynnikov:2006Monotonic), but both the representation of the knot as a grid diagram and the carrying out of the simplification steps are more complicated. These distinct representations of the same mathematical object lend themselves to different foundational models (Gukov:2021Learning; Kauffman:2022Rectangular); for example, braid word representations of knot projections are closer to natural language, whereas other representations are closer to vision tasks or graphs. The different representations also inform the type of algorithm that needs to be used. If the simplification is non-monotonic, a local search can get stuck in local minima.
@@
One common activity in mathematics, when confronted with a new statement of unknown truth, a conjecture, is to engage in the search for proof – or a counterexample. This comes with its own workflows. The search for counterexamples to conjectures is particularly useful for hard problems with a truth certificate that can be verified in polynomial time, such as NP-hard or NP-complete problems. The idea is to cast the search problem into a Markov Decision Problem (MDP) whose terminal states are counterexamples, and attempt to solve the MDP using data science techniques. In the past, deep reinforcement learning has proven to be a powerful tool. If the RL agent finds a solution to the MDP, its episodic rollouts provide truth certificates for the counterexample, thus establishing a verifiably correct proof by counter-example, see e.g. (Gukov:2024Rigor) for a recent summary of these ideas and (Gukov:2021Learning; Wagner:2021Constructing; Gukov:2023SearchingRibbons; Charton:2024Patternboost) for some recent concrete applications.
In the formal setting, Learning to Disprove frames counterexample search in Lean 4 as proposing a candidate counterexample and then certifying it by producing a Lean proof term, using a large synthetic corpus of counterexample instances (learningtodisprove2025).
4.2 Meta-Reasoning
Many of the current evaluations of mathematical reasoning abilities focus in reasoning inside a particular problem. In a mathematics competition or a class, that is generally enough: it is assumed that all problems are solvable, that solutions are not too long, and furthermore that all the necessary tools come from a known, bounded set (e.g., of topics allowed in competitions, or pre-requisites for the class). At the research frontier, however, a substantial amount of work is spent at the meta-level, where these assumptions are constantly being evaluated. The failure to prove a statement might be due to it being false (and perhaps a counter-example can be extremely non-trivial to produce). Or, the appropriate tools might not yet exist. In either case, the problem-level assumptions above might be violated, and one has to constantly decide whether to keep pushing on the current plan, and what to do if not. To the best of our knowledge, none of the current evaluations for mathematical reasoning directly probe these meta-level judgements, even though they represent an important opportunity for assistance from research copilots.
Some of the most basic meta-judgements that we might want to evaluate are (a) is a given statement true? (and does that change in light of given failed attempts to prove it?) (b) if it is, how much work will it take to prove it (or, given a partial proof, how much work is left?), and (c) is a given strategy adequate to prove it? In the formal setting, LeanProgress trains a predictor of proof progress from Lean proof states and uses it to guide search on longer Mathlib proofs (leanprogress2025). Many examples that we could use to evaluate the ability to answer these questions could be extracted from finished works: papers have examples of true statements; removing assumptions from those statements often renders them false (and most papers justify why the assumptions are needed, if not apparent, such as by constructing counter-examples where those assumptions are violated). We can measure the complexity of existing proofs in various ways (e.g., length, or by their dependencies inside a paper) to evaluate predictions of difficulty — these can also be made relative, e.g., identifying which lemma in the context of a broader proof is going to take the most work. On the other hand, many important meta-level decisions cannot be readily observed in finished work, such as when to change the problem locally (such as adding stronger assumptions), or switching problems entirely — perhaps temporarily, e.g., to gain intuition on a special case first, or more firmly, and in that case deciding what to work on instead. For such decisions, one might need to observe mathematicians doing their work, as opposed to their finished products, as we discuss next.
@@ 4.3 Data Collection in Real Environments
To gather representative data on mathematical workflow steps, as well as other metamathematical items, such as proof heuristics, limitations of certain proof techniques, etc., it may be desirable to observe and absorb the full process of producing mathematics, including all data that is not represented in written, textual form. For an example of real-time narration of thought processes that arise when doing competitive problem-solving, see co-author Tim Gowers’ series of YouTube videos. [^15] https://www.youtube.com/@TimothyGowers0/videos
@@
Unfortunately, many of these intermediate process steps become evident only implicitly from data sources. For example, at the elementary level, various workflows to solve various limits, such as $\lim_{x\to 0}\frac{\sin x-x}{x^{3}}$ or $\lim_{x\to\infty}x^{1/x}$, which require repeated applications of known theorems, such as de l’Hôpital’s theorem, or related ones to resolve indeterminate cases, are distilled only by solving a large number of exercises; no explicit annotations for these workflows exist. On a more advanced level, these are often conveyed in blogs, talks, and oral discussions between mathematicians at conferences – and not in the typical data sources used to train LLMs.
@@
In principle, one could implement longitudinal studies that track research projects from inception to completion. However, such data collection would require classifying, structuring, and recording different tasks throughout the mathematician’s daily work. The apparent challenge is to scale such approaches while keeping the overhead as small as possible for the researcher whose work is tracked. The KWARC group at FAU Erlangen-Nürnberg has developed infrastructure for semantic mathematical documents and tested such systems with hundreds of students per semester (kohlhase2006omdoc; berges2023learning), demonstrating the feasibility of structured mathematical data collection at scale, though their focus has been primarily on finished documents rather than the discovery process.
@@
While such approaches are welcome to be tested, as a first alternative step, we recommend transcribing from online sources such as lectures, panel discussions, seminars, vlogs, etc. VideoMathQA evaluates mathematical reasoning grounded in lecture-style videos, including long videos, with questions that require integrating visual and spoken information across time (videomathqa2025). While such data only requires minimal need for postprocessing and is often already collected in the form of videos, it typically covers broader mathematical aspects than textbooks. However, while humans can frequently generalize from a single instance of a certain workflow, LLMs often require a dataset that contains sufficiently many examples. While data collection effort will elicit higher-quality mathematical data, the time lag between the time when new research discoveries are made and when these are represented in lectures can be significant. On the other hand, at venues such as conferences, a lot of information about mathematics is exchanged, but not recorded. While it would technically be possible to record a portion of the conversations that happen at a conference in a privacy-preserving manner, and this would be a great source of data, as one would consistently have data that is at the forefront of research (unlike the mentioned panel discussions, where there is often a time lag), it would require a paradigm shift from conference participants to accept such data collecting measures. We urge further conversations about privacy-preserving ways of curating richer datasets on mathematical workflows, that minimally impinge on mathematicians’ naturalistic practices.
4.3.1 From Artifacts to Trajectories in Formal Proof Engineering
One distinctive advantage of formal mathematics is that the development environment already produces fine-grained signals that can be turned into datasets, which are in turn used to improve future systems. These signals include elaborator diagnostics, proof states, tactic traces, resource usage (heartbeats), and even lint warnings about unstable or redundant proof steps. This makes formal proof engineering a promising setting for process-level datasets.
One such dataset type is a repair trajectory corpus: sequences of (i) an attempted proof edit, (ii) compiler/LSP diagnostics (type errors, missing imports, failing goals, timeouts), and (iii) the subsequent patch. Such data directly represents the compile–diagnose–fix loop that dominates real formalization workflows, and it can be used to train copilots that are useful beyond one-shot “fill the sorry” tasks.
A second such dataset type is a refactoring corpus: pairs of proofs before/after refactoring steps that improve readability, reduce instance pollution, decrease resource usage, or align with library style. For example, proof refactoring guidance emphasizes extracting helper lemmas, compiling after each extraction, and using fast diagnostic feedback loops that (i) test compilation after each extraction, and then (ii) use LSP diagnostics to locate natural boundaries. Such datasets would support copilots that can shepherd code through review, not merely produce a valid proof.
As a third example, stability and hygiene linting (e.g., flagging tactic instability or unused steps) can be recorded as structured metadata. This enables benchmarks that measure not just whether a proof checks, but whether it is stable across versions and robust under realistic resource constraints.